OpenStudy (anonymous):

Consider a situation where immediately after birth, twins are separated. One continues to live on earth, while the other is whisked away in a space ship at 90% of the speed of light. After 65 years they are reunited. Will they look the same? Show proof of your answer through mathematics.

OpenStudy (schrodingers_cat):

Apply equation for time dilation \[T =\frac{ T _{o} }{ \sqrt{1 - \frac{v ^{2} }{ c^{2} }} }\]